a) Find a non-zero unit vector u with terminal point Q(-7,1,4) such that u has the opposite direction as v=(3,2,1).

b) Find a non-zero unit vector u with terminal point Q(-7,1,4) such that u has the same direction as v=(1,2,3).
1
Expert's answer
2020-03-09T13:59:10-0400

a) "\overrightarrow{PQ}=|\vec{u}|=1"

"Q(-7,1,4); P(x,y,z)"

"\vec{v}=(3,2,1)"

"\vec{u}\uparrow\downarrow\vec{v} \rArr \vec{u}=\lambda\vec{v}, \lambda<0"

"\begin{cases}
 |\vec{u}|=\sqrt{(-7-x)^2+(1-y)^2+(4-z)^2}=1 \\
 (-7-x,\ 1-y, \ 4-z)=\lambda(3,\ 2,\ 1)
\end{cases}
\hArr"

"\begin{cases}
 (-7-x)^2+(1-y)^2+(4-z)^2=1 \\
 -7-x=3\lambda\\
 1-y=2\lambda\\
 4-z=\lambda
\end{cases}
\hArr"

"\begin{cases}
 (-7-x)^2+(1-y)^2+(4-z)^2=1 \\
 (-7-x)^2=9\lambda^2\\
 (1-y)^2=4\lambda^2\\
 (4-z)^2=\lambda^2
\end{cases}
\hArr"

"\begin{cases}
 9\lambda^2+4\lambda^2+\lambda^2=1 \\
 (-7-x)^2=9\lambda^2\\
 (1-y)^2=4\lambda^2\\
 (4-z)^2=\lambda^2
\end{cases}
\hArr
\begin{cases}
 \lambda^2=\frac{1}{14} \\
 (-7-x)^2=\frac{9}{14}\\
 (1-y)^2=\frac{4}{14}\\
 (4-z)^2=\frac{1}{14}
\end{cases}
\hArr"

"\lambda<0 \rArr\lambda=-\frac{1}{\sqrt{14}}"

"\begin{cases}
 \lambda=-\frac{1}{\sqrt{14}} \\
 -7-x=-\frac{3}{\sqrt{14}}\\
 1-y=-\frac{2}{\sqrt{14}}\\
 4-z=-\frac{1}{\sqrt{14}}
\end{cases}
\hArr
\begin{cases}
 \lambda=-\frac{1}{\sqrt{14}} \\
 x=\frac{3}{\sqrt{14}}-7\\
 y=\frac{2}{\sqrt{14}}+1\\
 z=\frac{1}{\sqrt{14}}+4
\end{cases}
\hArr
\vec{u}=(-\frac{3}{\sqrt{14}},-\frac{2}{\sqrt{14}},-\frac{1}{\sqrt{14}})"


b) "\overrightarrow{PQ}=|\vec{u}|=1"

"Q(-7,1,4); P(x,y,z)"

"\vec{v}=(1,2,3)"

"\vec{u}\upuparrows\vec{v} \rArr \vec{u}=\lambda\vec{v}, \lambda>0"

"\begin{cases}
 |\vec{u}|=\sqrt{(-7-x)^2+(1-y)^2+(4-z)^2}=1 \\
 (-7-x,\ 1-y, \ 4-z)=\lambda(1,\ 2,\ 3)
\end{cases}
\hArr"

"\begin{cases}
 (-7-x)^2+(1-y)^2+(4-z)^2=1 \\
 -7-x=\lambda\\
 1-y=2\lambda\\
 4-z=3\lambda
\end{cases}
\hArr"

"\begin{cases}
 (-7-x)^2+(1-y)^2+(4-z)^2=1 \\
 (-7-x)^2=\lambda^2\\
 (1-y)^2=4\lambda^2\\
 (4-z)^2=9\lambda^2
\end{cases}
\hArr"

"\begin{cases}
 \lambda^2+4\lambda^2+9\lambda^2=1 \\
 (-7-x)^2=\lambda^2\\
 (1-y)^2=4\lambda^2\\
 (4-z)^2=9\lambda^2
\end{cases}
\hArr
\begin{cases}
 \lambda^2=\frac{1}{14} \\
 (-7-x)^2=\frac{1}{14}\\
 (1-y)^2=\frac{4}{14}\\
 (4-z)^2=\frac{9}{14}
\end{cases}
\hArr"

"\lambda>0 \rArr\lambda=\frac{1}{\sqrt{14}}"

"\begin{cases}
 \lambda=\frac{1}{\sqrt{14}} \\
 -7-x=\frac{1}{\sqrt{14}}\\
 1-y=\frac{2}{\sqrt{14}}\\
 4-z=\frac{3}{\sqrt{14}}
\end{cases}
\hArr
\begin{cases}
 \lambda=\frac{1}{\sqrt{14}} \\
 x=-\frac{1}{\sqrt{14}}-7\\
 y=-\frac{2}{\sqrt{14}}+1\\
 z=-\frac{3}{\sqrt{14}}+4
\end{cases}
\hArr
\vec{u}=(\frac{1}{\sqrt{14}},\frac{2}{\sqrt{14}},\frac{3}{\sqrt{14}})"